# Towards Robust Multimodal Sentiment Analysis with Incomplete Data

### v1

@@ Appendix A Implementation Details
We used PyTorch 2.2.1 to implement the method. The experiments were conducted on a PC with an AMD EPYC 7513 CPU at 2.6GHz and 1.0TB memory, and an NVIDIA Tesla A40 with 48GB memory. To ensure consistent and fair comparisons across all methods, we conducted each experiment three times using fixed random seeds of $1$, 1112, and 1113, respectively. Details of the key parameters used in the experiments are shown in Table 6.
In addition, the result of MISA, Self-MM, MMIM, CENET, TETFN, and TFR-Net is reproduced by the authors from open source code in the MMSA [^1] MMSA: https://github.com/thuiar/MMSA (Mao et al., 2022), which is a unified framework for MSA, using default hyper-parameters. The result of ALMT is reproduced by the authors from open source code on Github [^2] ALMT: https://github.com/Haoyu-ha/ALMT using default hyper-parameters.
Table 6: Hyper-parameters of LNLN we use on the different datasets
@@
Appendix B Additional Experiments and Analysis
B.1 Effect of Regularization Weight on Model Performance
@@
As shown in Table 7, we empirically tried different combinations of hyperparameters in the loss function. The results shows that the selected parameters can make the LNLN achieve the better performance in most metrics. This also demonstrates the effectiveness of the optimization objectives of LNLN.
@@
B.2 Analysis of Model Stability
To evaluate the stability of the model, we selected the MOSI dataset for ablation experiments. Specifically, based on the overall performance in Table 1, we selected several representative methods to additionally compute the standard deviation. For each method, we first calculate the standard deviation of evaluation results for the three random seeds when the missing rate $r$ ranges from 0 to 0.9, and then average the 10 sets of standard deviation. The final result is shown in Table 8. We can see that the standard deviation of these methods is not very large in most metrics, indicating that all these methods can guarantee stability in the presence of random noise. It should be noted that LNLN strikes a balance between overall performance and stability.
Table 8: Comparison of Model Stability on MOSI dataset. Note: -$\pm$- means mean±$\pm$±std. The smaller MAE indicate the better performance.
@@ B.2 Analysis of Model Stability
B.3 Analysis of Data Distribution
@@
The unbalanced data distribution makes it difficult for the model to perform well in data missing scenarios. Specifically, it is easy for the model to engage in lazy behavior in high-noise scenarios, simply biasing predictions towards categories with a higher proportion in the training set. These unbalanced distributions may further increase the learning difficulty of the model when facing data missing. More details can be found in Appendix B.5.
B.4 Details of Robust Comparison
@@
In addition, when $r$ is high, many models converge in their performance on metrics such as Acc-7 and Acc-5 (see Appendix B.5). This is due to the fact that these models exhibit lazy behavior, tending to predict inputs as categories with a higher number of samples in the training set. In such cases, we believe that the models do not actually learn effective knowledge, even if their performance appears good.
Figure 4: Data distribution of MOSI, MOSEI and SIMS datasets.
Table 9: Details of Robust Comparison on MOSI and MOSEI with different random missing rates. Note: The smaller MAE indicate the better performance. Note: The smaller MAE indicate the better performance.
@@
Table 10: Details of Robust Comparison on SIMS with different random missing rates.
@@
B.5 Visualization of Confusion Matrix
@@
Figure 7: Five-category Confusion matrix of several representative methods on SIMS dataset. Note: 0-4 denote weakly negative, negative, neutral, weakly positive, positive, respectively.
B.6 Visualization of Representations
Figure 8 shows the language feature $H^{1}_{l}$ extracted without applying data missing and the corrected dominant feature $H^{1}_{d}$ with different data missing rates. we can see that the model tends to learn $H^{1}_{d}$ that is consistent with the distribution of the $H^{1}_{l}$, indicating that LNLN can complement the information of the language modality when data is missing. Additionally, when the missing rate $r$ is 1.0 (no valid information for the input), it becomes difficult for the model to learn $H^{1}_{d}$ that is consistent with the language feature. This also suggests that the generator is indeed trying to learn $H^{1}_{d}$ for fusion in DMML, rather than just guessing.
@@ B.6 Visualization of Representations
B.7 Analysis of the generalization to modality missing scenarios
@@
Table 11: Generalization comparison of the overall performance on MOSI and MOSEI with random modality missing. Note: the parameters used for evaluation are consistent with those used for testing in random data missing. The smaller MAE indicate the better performance.
@@
Table 13: Generalization comparison on MOSI and MOSEI with random modality missing. Note: The smaller MAE indicate the better performance.
@@
Table 14: Generalization comparison on SIMS with random modality missing. Note: The smaller MAE indicate the better performance.
@@
Audio & Viusal Modalities Missing
@@
Appendix C Social Impacts
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope. They clearly outline the capabilities of the proposed LNLN and its performance in handling data missing scenarios, as well as emphasizing the comprehensive evaluation of existing methods and their significance in improving robust Multimodal Sentiment Analysis.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discussed the limitations of the work, which is specifically in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The article does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper reports the network structure required for reproduction as well as the hyperparameter settings. The code will be made public on Github after being accepted.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The datasets used in the thesis are open datasets and can be accessed by anyone upon request. In addition the code will be open-sourced at Github later.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Optimizer selection, data partitioning, and other details are described. Hyperparameters are chosen empirically.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: As shown in 8, we report the variance/fluctuation of several representative methods on the mosi dataset. It shows the overall stability of the different methods in the face of noise impacts of different levels.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report this information in Section Implementation Details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This study meets the requirements of moral and ethical norms.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We describe this in the appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The used data has been cited in the paper. The code will be available on Github after sort out.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: The dataset used is a public dataset, and the code will be available on Github after sort out.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve above problems.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This study does not involve the above issues and the datasets used are all publicly available datasets.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@ 4.3 Implementation Details
We used PyTorch 2.2.1 to implement the method. The experiments were conducted on a PC with an AMD EPYC 7513 CPU and an NVIDIA Tesla A40. To ensure consistent and fair comparisons across all methods, we conducted each experiment three times using fixed random seeds of $1$, 1112, and 1113. Details of the hyperparameters are shown in Table 1. In addition, the result of MISA, Self-MM, MMIM, CENET, TETFN, and TFR-Net is reproduced by the authors from open source code in the MMSA [^2] MMSA: https://github.com/thuiar/MMSA (Mao et al., 2022), which is a unified framework for MSA, using default hyperparameters. The result of ALMT is reproduced by the authors from open source code on Github [^3] ALMT: https://github.com/Haoyu-ha/ALMT.
Table 1: Hyperparameters of LNLN we use on the different datasets
@@
4.4 Robustness Comparison
Tables 2 and 3 show the robustness evaluation results on the MOSI, MOSEI, and SIMS datasets. As shown in Table 2, LNLN achieves state-of-the-art performance on most metrics. On the MOSI dataset, LNLN achieved a relative improvement of 9.46% on Acc-7 compared to the sub-optimal result obtained by MMIM, demonstrating the robustness of LNLN in the face of different noise effects. However, on the MOSEI dataset, LNLN achieves only sub-optimal performance on metrics such as Acc-7 and Acc-5. After analyzing the data distribution (see Appendix A.3) and the confusion matrix (see Appendix A.6), we believe that this is because LNLN does not overly bias the neutral category, which has a disproportionate number of samples in noisy scenarios. As shown in Table 3, LNLN achieves an improvement in F1 on the SIMS dataset, with a relative improvement of 9.17% on F1 compared to the sub-optimal result obtained by ALMT. Similar to CENET’s performance on the MOSEI dataset, TETFN on the SIMS dataset has a tendency to predict inputs as weakly negative categories with high sample sizes, resulting in seemingly better performance on some metrics. Additionally, as shown in Figure 2, we present the performance curves of several advanced methods under varying missing rates. The results demonstrate that the proposed LNLN consistently outperforms others across most scenarios, showing its robustness under different missing rates.
Table 2: Robustness comparison of the overall performance on MOSI and MOSEI datasets. Note: The smaller MAE indicates the better performance.
Method MOSI MOSEI
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-7 Acc-5 Acc-2 F1 MAE Corr
Table 3: Robustness comparison of the overall performance on SIMS dataset. Note: The smaller MAE indicates the better performance.
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
MISA 31.53 56.87 72.71 66.30 0.539 0.348
Self-MM 32.28 56.75 72.81 68.43 0.508 0.376
MMIM 31.81 52.76 69.86 66.21 0.544 0.339
CENET 22.29 53.17 68.13 57.90 0.589 0.107
TETFN 33.42 56.91 73.58 68.67 0.505 0.387
TFR-Net 26.52 52.89 68.13 58.70 0.661 0.169
ALMT 20.00 45.36 69.66 72.76 0.561 0.364
LNLN 34.64 57.14 72.73 79.43 0.514 0.397
Figure 2: Performance curves of various missing rates. (a), (b) and (c) are the F1 curves on MOSI, MOSEI, and SIMS, respectively. (d), (e) and (f) are the MAE curves on MOSI, MOSEI, and SIMS, respectively. Note: The smaller MAE indicates the better performance.
4.5 Effects of Different Components
To verify the role of different components, we present the results after subtracting each component on the MOSI dataset separately. As shown in Table 4, some metrics show an upward trend in performance when certain modules are removed. On one hand, we believe this is due to randomness in noisy scenarios. On the other hand, the same hyperparameters cannot make the model perform optimally across all metrics. Moreover, the performance of LNLN decreases relatively slightly when the DMC and reconstructor are removed individually. However, there is a significant drop in performance when both are removed, e.g., 4.72% decrease in Acc-7 and 5.87% decrease in Acc-5. Performance drops further after replacing DMML with concatenation fusion, which also proves that DMML plays an important role in MSA. Additionally, after removing the noisy data used for training, the performance of LNLN also shows a significant decrease. This observation indicates that LNLN needs to learn to be aware of noise with the help of noisy data. Overall, both the presence of noisy data and well-designed components are crucial for robust MSA in scenarios with random missing data.
Table 4: Effects of different components. Note: The smaller MAE indicates the better performance.
Method MOSI SIMS
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-5 Acc-3 Acc-2 F1 MAE Corr
4.6 Effects of Different Regularization
As shown in Table 5, we present the ablation results after subtracting each regularization on the MOSI dataset separately. Obviously, after each regularization is removed, LNLN shows a decrease in performance on most metrics. This demonstrates that regularization contributes to the learning of the model. Similar to the results in Table 4, there are a few metrics that improve after the removal of regularization. We believe this is mainly due to the difficulty of optimizing a single set of hyperparameters to achieve the best performance across all metrics.
Method MOSI SIMS
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-5 Acc-3 Acc-2 F1 MAE Corr
4.7 Effects of Different Modalities
To verify the role of different modalities, we present the ablation results after subtracting each modality from the MOSI dataset. As shown in Table 6, all modalities contribute to the final performance. Notably, when the language modality is removed, the performance of LNLN shows a significant decrease, while removing the visual and audio modalities results in a smaller decrease. This indicates that the language modality plays a crucial role in the MSA dataset. Moreover, we observe that LNLN converges on most metrics even when the visual and audio modalities are removed. This is due to the model’s lazy behavior, where LNLN tends to fix its predictions to a certain category. More details can be found in Appendix A.3 and A.6.
Table 6: Effects of different modalities. Note: The smaller MAE indicates the better performance.
Method MOSI
Acc-7 Acc-5 Acc-2 F1 MAE Corr
LNLN 34.26 38.27 72.55 / 70.94 72.73 / 71.25 1.046 0.527
w/o language 16.30 16.31 54.88 / 55.79 60.33 / 63.10 1.399 0.033
w/o audio 33.51 38.10 72.31 / 70.54 73.67 / 71.10 1.064 0.535
w/o vision 33.53 38.15 72.32 / 70.58 73.68 / 71.14 1.064 0.536
4.8 Discussion
In this work, we conduct a comprehensive evaluation of current advanced methods, the results of which have triggered some thoughts. Specifically, we find that all methods fail to accurately predict the inputs in high missing rate scenarios because most of them are designed specifically for complete data (see Appendix A.6 for more details). However, some methods show relatively more robust performance in low missing rate scenarios and some specific scenarios (see Appendix A.8), such as Self-MM, TETFN, and ALMT. We believe that some of the ideas in these methods may be useful for future research on more robust MSA for incomplete data.
Additionally, we explore the performance of these methods in modality missing scenarios, which is a special case of random data missing where the partial modality missing rate $r=1.0$. We find that when language modality is missing, the performance of many models decreases significantly and converges to the same value. More details can be seen in Appendix A.8.
5 Conclusion and Future Work
In this paper, a novel Language-dominated Noise-resistant Learning Network (LNLN) is proposed for robust MSA. Due to the collaboration between Dominate Modality Correction (DMC) module, the Dominant Modality based Multimodal Learning (DMML) module, and the Reconstructor, LNLN can enhance the dominant modality to achieve superior performance in data missing scenarios with varying levels. Extensive evaluation shows that none of the existing methods can effectively model the data with high missing rates. The related analyses can provide suggestions for other researchers to better handle robust MSA. In the future, we will focus on improving the generalization of the model to handle different types of scenes and varying intensities of noise effectively.
Limitations
We believe there are several limitations to the LNLN. 1) The LNLN achieves good performance in data missing scenarios, but is not always better than all other methods in the modality missing scenarios, which demonstrates the lack of multi-scene generalization capability of the LNLN. 2) The data in real-world scenarios is much more complex. In addition to the presence of missing data, other factors need to be considered, such as diverse cultural contexts, varying user behavior patterns, and the influence of platform-specific features on the data. These factors can introduce additional noise and variability, which may require further model adaptation and tuning to handle effectively. 3) Tuning the hyperparameters, particularly those related to the loss functions, can be challenging and may require more sophisticated methods to achieve optimal performance.
Acknowledgements
This work is supported by the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001).
Appendix A Additional Experiments and Analysis
A.1 Effect of Regularization Weight on Model Performance
@@
As shown in Table 7, we empirically tried different combinations of hyperparameters in the loss function. The results show that the selected parameters can make the LNLN achieve better performance in most metrics. This also demonstrates the effectiveness of the optimization objectives of LNLN.
@@
A.2 Analysis of Model Stability
To evaluate the stability of the model, we selected the MOSI dataset for ablation experiments. Specifically, based on the overall performance (mean) in Table 2, we selected several representative methods to additionally compute the standard deviation. For each method, we first calculate the standard deviation of evaluation results for the three random seeds when the missing rate $r$ ranges from 0 to 0.9, and then average the 10 sets of standard deviation. The final result is shown in Table 8. We can see that the standard deviation of these methods is not very large in most metrics, indicating that all these methods can guarantee stability in the presence of random noise. It should be noted that LNLN strikes a balance between overall performance and stability.
Table 8: Comparison of model stability on MOSI dataset. Note: -$\pm$- means mean±$\pm$±std. The smaller MAE indicates better performance.
@@ A.2 Analysis of Model Stability
A.3 Analysis of Data Distribution
@@
The unbalanced data distribution makes it difficult for the model to perform well in data missing scenarios. Specifically, it is easy for the model to engage in lazy behavior in high-noise scenarios, simply biasing predictions towards categories with a higher proportion in the training set. These unbalanced distributions may further increase the learning difficulty of the model when facing data missing. More details can be found in Appendix A.6.
A.4 Case Study
As shown in Figure 3, we visualize several successful and failed predictions made by LNLN and ALMT from the MOSI dataset for the case study. It shows that LNLN can perceive sentiment cues in challenging samples, which demonstrates its ability to capture sentiment cues in noisy scenarios. However, for inputs with high missing rates (e.g., the third example in the figure), both LNLN and ALMT fail to make correct predictions. We believe this is due to the high loss of valid information in the multimodal input, making accurate predictions difficult.
Figure 3: Visualization of successful and failed predictions. Note: The input is visualized to facilitate readers’ understanding. In practice, random data missing is applied to the original input sequence, as described in Section 3.
A.5 Details of Robust Comparison
@@
In addition, when $r$ is high, many models converge in their performance on metrics such as Acc-7 and Acc-5 (see Appendix A.6). This is due to the fact that these models exhibit lazy behavior, tending to predict inputs as categories with a higher number of samples in the training set. In such cases, we believe that the models do not actually learn effective knowledge, even if their performance appears good.
Figure 4: Data distribution of MOSI, MOSEI, and SIMS datasets.
Table 9: Details of robust comparison on MOSI and MOSEI with different random missing rates. Note: The smaller MAE indicates the better performance.
@@
Table 10: Details of robust comparison on SIMS with different random missing rates. Note: The smaller MAE indicates the better performance.
@@
A.6 Visualization of Confusion Matrix
@@
Figure 7: Five-category confusion matrix of several representative methods on SIMS dataset. Note: 0-4 denote weakly negative, negative, neutral, weakly positive, and positive, respectively.
A.7 Visualization of Representations
Figure 8 shows the language feature $H^{1}_{l}$ extracted without applying data missing and the corrected dominant feature $H^{1}_{d}$ with different data missing rates. we can see that the model tends to learn $H^{1}_{d}$, which is consistent with the distribution of the $H^{1}_{l}$, indicating that LNLN can complement the information of the language modality when data is missing. Additionally, when the missing rate $r$ is 1.0 (no valid information for the input), it becomes difficult for the model to learn $H^{1}_{d}$ that is consistent with the language feature. This also suggests that the generator is indeed trying to learn $H^{1}_{d}$ for fusion in DMML, rather than just guessing.
@@ A.7 Visualization of Representations
A.8 Analysis of the generalization to modality missing scenarios
@@
Table 11: Generalization comparison of the overall performance on MOSI and MOSEI with random modality missing. Note: the parameters used for evaluation are consistent with those used for testing in random data missing. The smaller MAE indicates better performance.
@@
Table 13: Generalization comparison on MOSI and MOSEI with random modality missing. Note: The smaller MAE indicates the better performance.
@@
Table 14: Generalization comparison on SIMS with random modality missing. Note: The smaller MAE indicates the better performance.
@@
Audio & Visual Modalities Missing
@@
Appendix B Social Impacts